Phc manual

for version 1.2 of Phc. Latest update: 12 July 2006.

Jia Wang (skyroam@gmail.com)

This manual documents version 1.2 of Phc, the validator of free proxy. The latest update of this manual: 12 July 2006.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005 Jia Wang. Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.2 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled "GNU Free Documentation License".

Table of Contents

1 What is phc? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Featurelist. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Requirements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4 Quick start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5 Introduction to basic configuration. . . . . . . . . . 6 6 Advanced configuration . . . . . . . . . . . . . . . . . . . . 9 7 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 8 Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 9 Format of proxylist . . . . . . . . . . . . . . . . . . . . . . . 12 10 Format of output . . . . . . . . . . . . . . . . . . . . . . . . 13 11 TODO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 GNU GENERAL PUBLIC LICENSE . . . . . . . . . 16

Preamble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND

MODIFICATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Appendix: How to Apply These Terms to Your New Programs . . . . . 21

Appendix A Copying This Manual . . . . . . . . . . 22

A.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . 22

A.1.1 ADDENDUM: How to use this License for your documents

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

12 Contact . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

1 What is phc?

Phc is a customizable proxy hunter. (It is not a really brave hunter, but a lazy and kind one who will only select beasts puted before him or her,:-).)

It reads free proxy list from file and outputs data used to value speed or response of the free proxies available. All of the process can be customized by the user.

Certainly, you can learn and finish configuration in a few minutes. And then, if you have a big proxylist, just have a cup of coffe to wait for the result.

2 Featurelist

Featurelist

1. Full commandline options. All of configuration can be controlled by commandline options which makes it feasible to customize phc from scripts.

2. Checking free http proxy^{1} with GET or CONNECT method:

• HTTP CONNECT

− This kind of http proxy accepts CONNECT request. Most of application based on tcp can use this kind of proxies, for example: telnet,ssh, http etc. − This kind of http proxy usually do not limit strictly the range of target port

of the request.

• HTTP GET

− The common http proxy. It only proxy common service, such as http,ftp,gopher. It is due to the proxy server what kind of service can be supported.

− Tt will benefit from cache of proxy server. So it is a good selection if many

people use the same proxy to visit similar sites in the same time.

3. Checking free socks5 proxy with connect method:

• SOCKS5 CONNECT

− Socks5 proxies accept connect command. This method of socks5 proxy is very like the connect method of http proxy. Just above, it can be used as tcp tunnel. That is to say, it can proxy any request based on TCP.

4. Checking socks4 proxy with connect method:

• SOCKS4 CONNECT

− Just above, this type of socks4 proxies can be used as tcp tunnel.

5. Checking the testproxy^{2} via myproxy.

Myproxy is the proxy which phc connect to it and go to testproxy via it. Now myproxy can be http (support CONNECT method) or socks5 (support connect command) proxy. (The support to socks4 myproxy in phc haven’t been open now.) Two kind of authentication of myproxy is supported for socks5 or http proxy: Notice, the directly connection in phc is open. So you can validate testproxy without firewall now.

1. Authentication with username and password needed. http: basic authentication socks5: username/password authentication(RFC1929).

2. no authentication.

That is to say, myproxy can be free proxy also. Certainly, you must have route to connect to myproxy directly.

6. Multithread.

1 proxies without authentication and free to use by the others 2 The free proxis in proxy list file

7. All delay on process are written to file and printed on screen in the same time. There is a simple script available used to sort the result according to the speed of proxies.

8. You can use ctrl-c to kill it without loss on result.

9. Read testproxy by line but not reading to memory at a glance. Phc is expected to check big proxylist without influencing the other process.

10. Recv timeout or send timeout can be set to avoid wasting time on slow proxies.

11. Either the file phc read proxylist from or the file phc output result can be customized.

12. Flexible, customizable. All detail on validating can be customized,such as: user agent: you can check testproxy as any browsers with it if you like. checksite: you can select the best testproxy for a single site. key: you can select a better key in the site of target to match the result thoroughly. ...

3 Requirements

Requirements

• CPU i386,alpha,sparc and more

• OS It has been tested on many linux dictribution, such as gentoo,debian ,fedora core. It works on SunOS and OSF1,too. I am not sure whether it will run at other operating system.

• Compile Libpthread is needed. It was implemented in most of the glibc in the world. I compile it sucessfuly with gcc. I am not sure whether it will be compiled by the other compilers.

• Network A http or socks5 proxy which support CONNECT. This type of firewall is described as myproxy just as above. Direct connection to testproxy without myproxy is supported without well test. Just let mytype=-1 to enable it. Marvin have written a patch for phc-1.0 to support connection directly without myproxy. If you met problems with mytype=-1 , you can also try this patch.

• Privilege. Common user who can create,read,write socket and file is ok. If you are puzzled ,I am sure you have this privilege, :-). Just try it.

4 Quick start

You can use phc after a few minutes if you do like this step by step:

1. Extract from source and enter the work dir:

     tar zxvf phc-1.2.tar.gz
     cd phc-1.2

2. Compile && Install:

     ./configure
     make
     make install

3. Phc will search configureation file and proxylist file in current dir. If you still in work

dir, just create configuration file like this:

     cp doc/phc.conf.example phc.conf

You can put phc.conf to /etc/phc.conf or ~/.phc.conf. Phc will search ./phc.conf,then ~/.phc.conf , at last /etc/phc.conf until it found the first one. Now read the configuration file "phc.conf". Notice: The line begin with ’#’ is commet line. The line after the line begin with ’#’ must be customized with you own setting. And the line begin with ’#! if’ need be updated only if the condition expression after "if" is true.

5 Introduction to basic configuration

• mytype The type of myproxy, which phc connect to directly. 1 for http proxy without authentication. 2 for socks5 proxy without authentication. 3 for http proxy with basic authentication. 4 for socks5 proxy with username/password authentication. -1 for connection to testproxy without myproxy(Testing).

• myhost The ip of domain name of myproxy. This will be in effect only if mytype != -1

• myport The http proxy service port of myproxy. This will be in effect only if mytype !=-1

• myuser This will be read only if mytype=3 or 4.

• mypass This will be read only if mytype=3 or 4.

For example, you username at myproxy is abc, password is guest, then you should set myuser and mypass like this:

     myuser=abc
     mypass=guest

• httptestmethod The http method which should be supported by the testproxy. httptestmethod=0(HTTP GET) or 1(HTTP CONNECT) is supported now. If you need a HTTP CONNECT proxy, just set it to 1: httptestmethod=1

• socks5testmethod The socks5 command which should be supported by the testproxy. Only 2 support to (SOCKS5 CONNECT) is implemented now. You should leave it intact now: socks5testmethod=2

• socks4testmethod BIND isnt’t implemented for socks4. So you should leave it intact: socks4testmethod=5

• logfilename The logfile with full path. If you prefer screen output:

logfilename=. If you want to see it laterly: logfilename=./phc.log

4. If you have proxy list file like the file proxylist in directory doc, just rename it as proxylist. Otherwise, you can copy the proxylist file in doc to current directory to begin test immediately:

 cp doc/proxylist .

Notice, phc.conf must be in the same directory now.

5. At last, start checking:

 phc

Tips: If you have found enough free proxies, you can use "^C" to stop phc at any time. The result of validating will be printed on screen and saved to file named proxyout in current directory immediately after any free proxy is found .

6. Example of configuration.

• Conditions: a socks5 proxy(socks.you.com) with username/password authentication. socks5 service port:1080, username:abc,password:def. Want to validate: free http CONNECT proxy and free socks5 proxy which support connect command. You configuration should look like this:

#You must customize the follow
mytype=4
myhost=socks.you.com
myport=1080
myuser=abc
mypass=def
httptestmethod=1
#You should leave the following items intact.
socks5testmethod=2
socks4testmethod=5
#Optional to be changed. But you SHOULD NOT comment or remove it.
threads=50
checksite=http://www.google.com:80/webhp?hl=en
key=2006 Google
user_agent=User-Agent: MSIE 6.0
myconreplysize=256
testconreplysize=256
testtagreplysize=65535
rcvtimeo=10
sndtimeo=5
logfilename=phc.log

• Conditions: a http proxy(proxy.you.com) without authentication. http proxy service port:8080. Want to validate: Common free http proxy and free socks5 proxy which support connect command. Your setting will be similiar with this:

#The follow lines must be customized.
mytype=1
myhost=proxy.you.com
myport=8080
httptestmethod=0
#Please leave the following items intact.
socks5testmethod=2
socks4testmethod=5
#The follow lines is optional to be changed , but you should not
#delete or remove it.
threads=50
checksite=http://www.google.com:80/webhp?hl=en
key=2006 Google
user_agent=User-Agent: MSIE 6.0

myconreplysize=256
testconreplysize=256
testtagreplysize=65535
rcvtimeo=10
sndtimeo=5
logfilename=phc.log

6 Advanced configuration

Please read See Chapter 4 [Quick start], page 5. to know the meaning of ’#!’ or ’#! if’ at the start of a line and what are the vars that must be customized.

Synatx: a line will be ignored unless it is in the format like this:

 varname=value

Notice: No chars (such as whitespace,quote) will be exceptional. So any chars before, after or inside "varname=value" will be recognized as part of "varname=value".

The list of advanced vars:

• threads The number of threads opened at the same time. This should be set according to ulimit. The number can’t be more than 65535. Though nocache header is used, it is expected that less threads give more precise delay information to evaluate the proxies better.

• myconreplysize The size of buffer used to receive reply after sending CONNECT request for testproxy to myproxy. This is used for some weird proxies. Please leave it intact if unsure.

• testconreplysize The size of buffer used to receive reply after sending CONNECT request for target to testproxy. It is only be used when you validate HTTP CONNECT proxies. Just as above, do not change it if unsure.

• testtagreplysize The size of buffer used to receive reply after sending GET request to the target. Phc will find key from this buffer.

• user agent Identified as Browsers or others to "cheat" some proxies.

• rcvtimeo Receiving timeouts in second.

• sndtimeo Sending timeouts in second.

• logfilename Redirect stderr to log file. Use a single "." ("." is just before LF) to leave stderr as it is.

• phcin The input file with full path which phc read proxylist from. If you do not set it, phc will use file named "proxylist" in the current directory.

• phcout The output file with full path which phc output result to. If you do not set it, phc will user file nameed "proxyout" in the current directory.If this file do not exists, phc will creat it. Otherwise, it will be overwrote.

The line begin with other strings which do not appeare here or See Chapter 4 [Quick start], page 5. will be ignored.

7 Options

Invoke it as: ./phc –help

You can see the table of map from commandline options to configuration options like this: ./phc: invalid option – - Phc 1.2, a customizable proxy hunter. Usage: phc [OPTION]...

-V display the version of Phc and exit.

All of the other options need arguments!

Commandline only options: -f the argument is configuration file

Configuration options: -t threads -g user agent -m mytype, the type of myproxy. -p myport, the port of myproxy. -h myhost, the ip or domainname of myproxy. -u myuser, the username of myproxy if need. -s mypass, the password of myproxy if need. -c checksite, the target site used to check testproxies. -k key, the keyword used to check whether testproxies give correct response. -r rcvtimeo, socket recv timeout in seconds. -n sndtimeo, socket send timeout in seconds. -S myconreplysize -T testconreplysize -G testtagreplysize -H httptestmethod -5 socks5testmethod -4 socks4testmethod -l logfilename -i phcin, the proxylist file used as input of phc. -o phcout, the file which phc output result to.

Notice: The option ‘-f’ is used to customize configuration file which haven’t been used in configuration file. The option ‘-V’ is used to print version of phc.

8 Scripts

There are two scripts in directory scripts.

• creatlist A script used to filter and create proxylist file from raw file. Only line in format of "ip:port@proto.*" will be reserved and the chars after proto in the line will be removed. All duplicate line will be removed. What the script does is: start->read aline from raw file->capitalize ->find port between : and @->port is number->check the string after @ to find proto -> proto is SOCKS5 or HTTP -> write ip:port@proto to tmpfile -> continue read next line ... loop If any of the steps failed, just break and continue with the next line from start. After all of lines are filtered,changed and saved to tmpfile, sort and uniq are used to remove duplicate line.

Usage: ./createlist infile outfile Example:
     ./createlist p proxylist

• sortlist

sort the file proxylist. Usage:

     cp doc/sortlist .
     ./sortlist

9 Format of proxylist

The format of proxylist accepted by phc should like this: ip:port@proto.*

or ip:port

I remove support on domain, so ip must be IPV4 address. The port must be integer range from 1 to 65535. The proto can only be HTTP,http,socks5,SOCKS5,socks4 or SOCKS4. All of the chars afterproto will be ignored.

If no @ in the line, proto is default as HTTP. Differently, no other chars after port can be appeared.

Example: 61.132.89.130:80@HTTP 202.123.52.156:8080

10 Format of output

The current version of phc output timeouts of every steps of validation:

Method  Output (time increase ->)
HTTP GET  host:port@HTTP estab pget end
HTTP CONNECT  host:port@HTTPC estab pcon get end
SOCKS5 CONNECT  host:port@SOCKS5 estab pcon get end

The figure on process of validation. See the table for the meaning of symbol in the figure:

symbol  meaning
–  start to send, the left string is time.(such as ta)
<-  The data entered,the left string is the time when data come in.

Figure:

Phc – (myproxy) — testproxy – target(checksite) ta —————->

estab = tb - ta tb <—————- tc ——————————>

pcon = td - tc HTTP CONNECT ,SOCKS5 CONNECT or SOCKS4 CONNECT. td <—————————— te ——————————>

(p)get = tf - te get:HTTP CONNECT,SOCKS5 CONNECT,SOCKS4 CONNECT,

pget:HTTP GET

tf <—————————— tg ——————————>

end = th - tg th <——————————

delay  define

estab Delay from the time(ta) phc send CONNECT request for testproxy to myproxy (or connect to testproxy directly) to the time(tb) it receives "Connection Established".

pcon Delay from the time(tc) phc send CONNECT request for target to testproxy to the time(td) it receives "Connection Established". With HTTP GET method, it is impossilbe to get "connection delay" for that is added to pget delay. It is only be used with HTTP CONNECT or SOCKS5 CONNECT method.

get Delay from the time(te) phc send HTTP GET request to target to the time(tf) it receives the first byte data returned. This is only be used with HTTP GET method and the "middle-delay"(like pcon) can’t be got by the client. pget Delay from the time(te) phc send GET request for target to testproxy to the time(tf) it receives the first byte data returned.

end Delay from the time(tg or tf) phc receive the first byte data to the time time(th) it receives data including key.

The value of the delays: estab resp of TP pcon resp of TP + relay of TP + resp of TG get resp of TP + relay of TP + resp of TG + servicestart of TG pget 2*resp of TP + 2*relay of TP + 2*resp of TG + servicestart of TG end 0.5*resp of TP + 0.5*relay of TP + 0.5*resp of TG + serviceend of TG

The guide to use the delay

• If you want to select a testproxy with best response to myproxy(or phc), just sort the result according to estab.

• If you need a testproxy through which you can get the best response (shortest route?) from the target defined in configuration, please sort according to pcon , get or pget.

• On the other hand , if you care for the bandwith of the whole route to target, you should sort according to end.

Tips:

• You can sort testproxies according to estab if you want to browse webpages via testproxy(through myproxy firstly or not). This is the common choice if you need quick response ,dynamic targets to access and little data to transfer.

• You can sort testproxies according to end if you want to download files via testproxy(through myproxy firstly or not) from target. Bandwith but not delay is important to bulk download.

• The less delay, the more speed. You should select the testproxies with the shortest delay, such as estab,pcon,get,pget or end.

11 TODO

• SOCKS5 UDP

• https.

• cgi

• text interface?

• graphic interface? A very lame gtk interface to phc is developed by me.

• sort in checking. It may wast plenty of time. For large number of proxylist, I guess it will be awfull.

• switchproxy I have written a few scripts and simple codes to select best proxies. The main scripts update proxies as frequently as you like. more... updating Please contact with me if you have idears on phc.

GNU GENERAL PUBLIC LICENSE

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc.

51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

1. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. You must cause the modified files to carry prominent notices stating that you

changed the files and the date of any change. b. You must cause any work that you distribute or publish, that in whole or in part

contains or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

c. If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it.

Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program. In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following: a. Accompany it with the complete corresponding machine-readable source code,

which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable. If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

4. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

5. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

6. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

7. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program. If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances. It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

8. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

9. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

10. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

11. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PERMITTED BY APPLI-CABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPY-RIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MER-CHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

12. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS Appendix: How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

      one line to give the program's name and a brief idea of what it does.
      Copyright (C) yyyy  name of author
      This program is free software; you can redistribute it and/or modify
      it under the terms of the GNU General Public License as published by
      the Free Software Foundation; either version 2 of the License, or
      (at your option) any later version.
      This program is distributed in the hope that it will be useful,
      but WITHOUT ANY WARRANTY; without even the implied warranty of
      MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
      GNU General Public License for more details.
      You should have received a copy of the GNU General Public License
      along with this program; if not, write to the Free Software
      Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA  02110-1301, USA

Also add information on how to contact you by electronic and paper mail.

If the program is interactive, make it output a short notice like this when it starts in an interactive mode:

      Gnomovision version 69, Copyright (C) 19yy name of author
      Gnomovision comes with ABSOLUTELY NO WARRANTY; for details type `show w'.
      This is free software, and you are welcome to redistribute it
      under certain conditions; type `show c' for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

     Yoyodyne, Inc., hereby disclaims all copyright interest in the program
     `Gnomovision' (which makes passes at compilers) written by James Hacker.
     signature of Ty Coon, 1 April 1989
     Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

Appendix A Copying This Manual

A.1 GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

A.1.1 ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.2
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

12 Contact

• Reporting bugs: http://savannah.nongnu.org/bugs/?group=phc
• Homepage: http://www.nongnu.org/phc
• Online summary: http://savannah.nongnu.org/projects/phc
• Files: http://savannah.nongnu.org/files/?group=phc
• Email: skyroam@gmail.com

Index 30

Index

C

chapter,Advanced configuration . . . . . . . . . . . . . . . . 9 chapter,Featurelist . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 chapter,Format of output. . . . . . . . . . . . . . . . . . . . . 13 chapter,Format of proxylist . . . . . . . . . . . . . . . . . . . 12 chapter,Introduction to basic configuration. . . . . . 5 chapter,Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 chapter,Quick start . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 chapter,Requirements . . . . . . . . . . . . . . . . . . . . . . . . . 4 chapter,Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 chapter,TODO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

E

Example of configuration. . . . . . . . . . . . . . . . . . . . . . 6

F

FDL, GNU Free Documentation License . . . . . . . 22